OpenStudy (sagewilson):

Help please, will medal. A cone is inscribed in a regular square pyramid. If the pyramid has a base edge of 6" and a slant height of 9", find the volume of the cone. 12√2 18√2 27

OpenStudy (sagewilson):

@mathmale @zepdrix

OpenStudy (sagewilson):

I don't think its 27.

OpenStudy (thatonekidkobe):

http://openstudy.com/updates/4e1b7acb0b8bc22757480494 bruh

OpenStudy (sagewilson):

"your answer is 25.47 in^3" That's not even an answer choice.

OpenStudy (thatonekidkobe):

If by "find the volume of the cone," you mean find the volume of the cone inside the pyramid then ok. I don't know what you mean otherwise.

OpenStudy (sagewilson):

How would I go about putting it in square root form?

OpenStudy (thatonekidkobe):

Assuming the cone is inside the pyramid and the cone has a diamter of 6 aka a radius of 3 the slant height is 9 and with Pythagorean therom 9^2 = 3^2 + height^2 therefore height = sqrt 72 Area of cone is 1/3 * pi * r^2 * h 1/3 * pi * 3^2 * sqrt(72) 3*pi*sqrt(72) = about 79.972 inches cubed

OpenStudy (sagewilson):

72√2

OpenStudy (sagewilson):

The answer is 18√2π